Yoho Corp. issued $500,000 of its ten-year 6% bonds at 104. Each $1,000 bond carries ten warrants. Each warrant allows the holder to purchase one share of $10 par common stock for $50. Following the sale, the relevant market values were:

Bonds $980 (ex rights)
Warrants $14 each
Common Stock $60 each


The entry to record the sale of the bonds would include a

a. credit to Common Stock Warrants for $65,000.

b. debit to Discount on Bonds Payable for $10,000.

c. credit to Bonds Payable for $490,000.

d. debit to Cash for $560,000.
Answer #1

Convertible Bonds have features of both equity as well as liability. Since the convertible bonds have features of both liability (debt) as well as equity, it makes more sense to account for the liability portion and equity portion separately.

The split between equity and liability portion needs to be accounted for at the time of the issue of bonds itself. The equity & liability portion for the convertible bonds can be calculated using the Residual Approach. This approach assumes that the value of the equity portion is equal to the difference between the total amount received from the proceeds of the bonds and the present value of future cash flows from the bonds.

In the given question,

Fair Value of the bond (ex - right) = 980

Face value of the bond = 1000

Thus discount on issue of bond (ex right) = 1000 - 980 = 20

Issue price of the bond = 1000 * 104% = 1040

Total number of bonds issued = $500000/$1000 = 500 bonds

Therefore the journal entry to record the issuance of bond would be as follows:

Particulars Debit Credit
Cash 520000
Discount On Issue of bonds 10000
Bonds payable 500000
Warrant (plug) 30000

Thus the answer to given question would be:

b. debit to Discount on Bonds Payable for $10,000.

Kindly Up Vote